# Towards Understanding the Robustness of Diffusion-Based Purification: A Stochastic Perspective

### v2

@@ Abstract
Diffusion-Based Purification (DBP) has emerged as an effective defense mechanism against adversarial attacks. The efficacy of DBP has been attributed to the forward diffusion process, which narrows the distribution gap between clean and adversarial images through the addition of Gaussian noise. Although this explanation has some theoretical support, the significance of its contribution to robustness remains unclear. In this paper, we argue that the inherent stochasticity in the DBP process is the primary driver of its robustness. To explore this, we introduce a novel Deterministic White-Box (DW-box) evaluation protocol to assess robustness in the absence of stochasticity and to analyze the attack trajectories and loss landscapes. Our findings suggest that DBP models primarily leverage stochasticity to evade effective attack directions, and their ability to purify adversarial perturbations can be weak. To further enhance the robustness of DBP models, we introduce Adversarial Denoising Diffusion Training (ADDT), which incorporates classifier-guided adversarial perturbations into diffusion training, thereby strengthening the DBP models’ ability to purify adversarial perturbations. Additionally, we propose Rank-Based Gaussian Mapping (RBGM) to make perturbations more compatible with diffusion models. Experimental results validate the effectiveness of ADDT. In conclusion, our study suggests that future research on DBP can benefit from the perspective of decoupling the stochasticity-based and purification-based robustness.
@@ 1 Introduction
Figure 1: Comparison of attack trajectories under different evaluation settings. The attack trajectory in the standard White-box setting deviates significantly from the DW-box trajectory and shows lower effectiveness.
@@
Deep learning has achieved remarkable success in various domains, including computer vision (He et al., 2016), natural language processing (OpenAI, 2023), and speech recognition (Radford et al., 2022). However, in this flourishing landscape, the persistent specter of adversarial attacks casts a shadow over the reliability of these neural models. Adversarial attacks for a vision model involve injecting imperceptible perturbations into input images to trick models into producing false outputs with high confidence (Goodfellow et al., 2015; Szegedy et al., 2014). This inspires a large amount of research on adversarial defense (Zhang et al., 2019; Samangouei et al., 2018; Shafahi et al., 2019; Wang et al., 2023).
@@
Diffusion-based purification (DBP) (Nie et al., 2022) has recently gained recognition as a powerful defense mechanism against a range of adversarial attacks. Existing literature suggests that the robustness provided by DBP is primarily due to the forward diffusion process that narrows the distribution gap between clean and adversarial images through the application of Gaussian noise (Nie et al., 2022; Wang et al., 2022). However, although the reduction of the distribution gap is theoretically proven, its contribution to DBP robustness has not been sufficiently validated by empirical studies. Meanwhile, it is observed that the stochasticity of DBP may also contribute to the robustness (Nie et al., 2022).
@@
In light of this, we introduce an alternative perspective that highlights the role of stochasticity throughout the DBP process as a key contributor to its robustness, challenging the traditional focus on the forward diffusion process. To evaluate the impact of stochasticity, we employ a Deterministic White-box (DW-box) attack setting where the attacker has complete knowledge of both the model parameters and the stochastic elements. Our findings reveal that DBP models significantly lose their robustness when the process is entirely deterministic to the attacker, thereby emphasizing the critical importance of stochasticity. Further investigations into attack trajectories and the loss landscape demonstrate that DBP models do not counter adversarial perturbations by a flat loss landscape as adversarial training (AT) (Madry et al., 2018); instead, they rely on stochasticity to circumvent the most effective attack direction, as depicted in Figure 1.
@@
Building on our new perspective regarding DBP robustness, we hypothesize that it can be further enhanced by improving the capability of the diffusion model to purify adversarial perturbations. To test this hypothesis, we propose Adversarial Denoising Diffusion Training (ADDT) for DBP models. This method follows an iterative two-step process: first, the Classifier-Guided Perturbation Optimization (CGPO) step generates adversarial perturbations; then, the diffusion model training step updates the parameters of the diffusion model using these perturbations. To better integrate these perturbations within the diffusion framework, we introduce Rank-Based Gaussian Mapping (RBGM), which adjusts the adversarial perturbations to more closely resemble Gaussian noise, in alignment with the theory behind diffusion models. Experiments across various diffusion methods, attack settings, and datasets suggest that ADDT can consistently enhance DBP models’ robustness and purification ability. With further empirical analysis and discussions, we argue that future research on DBP should decouple the robustness based on stochasticity and that achieved by purification, which suggests two orthogonal directions for improving DBP: (1) enhancing its capability to purify adversarial perturbations with efficient training methods, and (2) defending Expectation of Transformation (EoT) attacks by increasing the variance of attack gradients.
@@
  • We present a novel perspective on DBP robustness, emphasizing the critical role of stochasticity and challenging the conventional purification-based belief that robustness primarily stems from reducing the distribution gap via the forward diffusion process.
  • We introduce a new Deterministic White-box attack scenario and show that DBP models depend on stochastic attack gradients to avoid the most effective attack directions, demonstrating distinct properties compared to robust models obtained by adversarial training.
  • Based on the proposed ADDT, we validate that the DBP robustness can be further enhanced by improving the capability of the diffusion model to purify adversarial perturbations.
@@ 2 Related Work
Adversarial training (AT). First introduced by Madry et al. (2018), AT seeks to develop a robust classifier by incorporating adversarial examples into the training process. It has nearly become the de facto standard for enhancing the adversarial robustness of neural networks (Gowal et al., 2020; Rebuffi et al., 2021; Athalye et al., 2018). Recent advances in AT harness the generative power of diffusion models to augment training data and prevent AT from overfitting (Gowal et al., 2021; Wang et al., 2023). However, the application of AT to DBP methods has not been thoroughly explored.
Adversarial purification. Adversarial purification utilizes generative models to remove adversarial perturbation from inputs before they are processed by downstream models. Traditionally, generative adversarial networks (GANs) (Samangouei et al., 2018) or autoregressive models (Song et al., 2018) are employed as the purifier model. More recently, diffusion models have been introduced for adversarial purification, in a technique termed diffusion-based purification (DBP), and have shown promising results (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2020a; Nie et al., 2022; Wang et al., 2022; Wu et al., 2022; Xiao et al., 2022). The robustness of DBP models is often attributed to the wash-out effect of Gaussian noise introduced during the forward diffusion process. Nie et al. (2022) propose that the forward process results in a reduction of the Kullback-Leibler (KL) divergence between the distributions of clean and adversarial images. Gao et al. (2022) suggest that while the forward diffusion process improves robustness by reducing model invariance, the backward process restores this invariance, thereby undermining robustness. However, these theories explaining the robustness of DBP models lack substantial experimental support.
@@ 3 Preliminaries
Adversarial training. Adversarial training aims to build a robust model by including adversarial samples during training (Madry et al., 2018). This approach can be formulated as a min-max problem, where it first generates adversarial samples (the maximization) and then adjusts the parameters to resist these adversarial samples (the minimization). Formally, this can be represented as:
${\min}_{\bm{\theta}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[{\max}_{\bm{\delta}\in B}L(f(\bm{\theta},\bm{x}+\bm{\delta}),y)\right],$ (1)
where $L$ is the loss function, $f$ is the classifier, $(\bm{x},y)\sim\mathcal{D}$ denotes sampling training data from distribution $\mathcal{D}$, and $B$ defines the set of permissible perturbation $\bm{\delta}$.
Diffusion models. Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) and Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) simulate a gradual transformation in which noise is added to images and then removed to restore the original image. The forward process can be represented as:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),$ (2)
where $\bm{x}_{0}$ is the original image and $\bm{x}_{t}$ is the noisy image. $\overline{\alpha}_{t}$ is the cumulative noise level at step $t$ ($1<t\leq T$, where $T$ is the number of diffusion training steps) . The model optimizes the parameters $\bm{\theta}$ by minimizing the distance between the actual and predicted noise:
$\bm{\theta}^{*}={\arg\min}_{\bm{\theta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},t)\|_{2}^{2}\right],$ (3)
where $\bm{\epsilon}_{\bm{\theta}}$ is the model’s noise prediction, with $\bm{\epsilon}_{\bm{\theta}}$, we can predict $\hat{\bm{x}}_{0}$ in a single step:
@@
where $\beta_{t}$ is the noise level at step $t$, $\hat{\bm{x}}_{t-1}$ is the recovered image in step $t-1$, $\bm{\epsilon}$ is sampled from $\mathcal{N}(0,\bm{I})$. DDIM proposes to speed up the denoising process by skipping certain intermediate steps. Recent work suggests that DDPM may also benefit from a similar approach (Nichol & Dhariwal, 2021). Score SDEs (Song et al., 2020b) give a score function view of DDPM and further lead to the derivations of DDPM++ (VPSDE) and EDM  (Karras et al., 2022). In this diffusion process, the noise terms $\bm{\epsilon}$ in equations  equation 2 and  equation 5 represent the key stochastic elements that govern the randomness of the process. These stochastic elements will be further elaborated in Section C.1.
@@ 4.1 Stochasticity as the Main Factor of DBP Robustness
As discussed in Section 2, previous studies primarily attribute the robustness of DBP to the forward diffusion process, which introduces Gaussian noise to both clean and adversarial images, thereby narrowing the distribution gap between them (Wang et al., 2022; Nie et al., 2022). As a result, adversarial perturbation can be “washed out” by Gaussian noise. However, it is also found that the robustness of DiffPure can be reduced by switching the SDE sampling to ODE, which introduces less randomness, implying the potential contribution of stochasticity to DBP robustness (Nie et al., 2022).
To assess whether stochasticity has a significant influence on DBP robustness, we implement DDPM and DDIM within the DiffPure framework (Nie et al., 2022), resulting in DPDDPM and DPDDIM, respectively. Note that DDPM is mathematically equivalent to the DDPM++ (VPSDE) model used in DiffPure but employs a smaller UNet architecture. DDIM, on the other hand, is a mathematical improvement over DDPM that facilitates an ODE-based diffusion process. DPDDPM introduces Gaussian noise in both the forward and reverse processes, making the entire process stochastic. In contrast, DPDDIM introduces Gaussian noise only in the forward process, and the reverse process is deterministic. The clean and robust accuracy of the two models on CIFAR-10 (Madry, 2017; Krizhevsky et al., 2009) under white-box PGD+EoT (Athalye et al., 2018) attack (as detailed in Section 6.1) are presented in Figure 2 (Clean and White). Although DPDDPM achieves higher clean accuracy, it exhibits lower robust accuracy under adaptive white-box attacks, consistent with the observation by Nie et al. (2022). However, this comparison is insufficient to reveal the full role of stochasticity in DBP robustness, as the forward processes of both DDPM and DDIM are stochastic.
@@
Figure 2: DPDDPM and DPDDIM robust accuracy under different attack settings on CIFAR-10. Both models lose most of their robustness only when the attacker knows all stochastic elements (DWBoth-box for DPDDPM and DWFwd-box for DPDDIM).
To isolate the impact of stochasticity, we introduce a new attack scenario called the Deterministic White-Box (DW-box) setting. In this setting, the attacker has full knowledge of not only the model parameters but also the specific sampled values for the stochastic elements used during evaluation, effectively rendering the diffusion process deterministic from the attacker’s perspective. This setting can be realistic if the attacker is aware of the seed or initial random state for the pseudo-random number generation utilized by the model. Concretely, we define three levels of attacker knowledge for our evaluations: (1) the conventional White-box setting, where the attacker has access to the model parameters but not the stochastic elements; (2) DWFwd-box/DWRev-box setting, where the attacker knows the stochastic elements in the forward/reverse process, in addition to the model parameters; (3) DWBoth-box setting, where the attacker has full knowledge of the model parameters and all the stochastic elements in both the forward and reverse processes. Details of these settings are provided in Section C.2.
We evaluated adversarial robustness on CIFAR-10 using $l_{\infty}$ attacks (see Section 6.1). Traditional theories emphasize forward diffusion as the primary defense mechanism, suggesting that both DPDDPM and DPDDIM should behave similarly under the DWFwd-box setting. However, if stochasticity throughout the diffusion process is crucial, DPDDIM, which becomes deterministic under the DWFwd-box setting, should experience a notable reduction in robustness, similar to DPDDPM in the DWBoth-box setting. As shown in Figure 2, in the DWFwd-box setting, DPDDPM maintains a significant portion of its robustness, whereas DPDDIM loses almost all of its resistance to adversarial attacks. Furthermore, DPDDPM exhibits a substantial drop in robustness only when the attacker has full access to both the forward and reverse stochastic elements, as seen in the DWBoth-box setting. This suggests that stochasticity across both the forward and reverse diffusion processes plays a critical role in maintaining robustness, challenging the conventional focus on forward diffusion alone.
Our findings suggest that DBP models primarily use stochasticity to resist adversarial attacks, rather than mainly depending on forward diffusion to mitigate adversarial perturbations, and it also reveals that DBP itself lacks the ability to effectively purify adversarial perturbations.

### v3

@@ Abstract
Diffusion-Based Purification (DBP) has emerged as an effective defense mechanism against adversarial attacks. The success of DBP is often attributed to the forward diffusion process, which reduces the distribution gap between clean and adversarial images by adding Gaussian noise. While this explanation is theoretically sound, the exact role of this mechanism in enhancing robustness remains unclear. In this paper, through empirical analysis, we propose that the intrinsic stochasticity in the DBP process is the primary factor driving robustness. To test this hypothesis, we introduce a novel Deterministic White-Box (DW-box) setting to assess robustness in the absence of stochasticity, and we analyze attack trajectories and loss landscapes. Our results suggest that DBP models primarily rely on stochasticity to avoid effective attack directions, while their ability to purify adversarial perturbations may be limited. To further enhance the robustness of DBP models, we propose Adversarial Denoising Diffusion Training (ADDT), which incorporates classifier-guided adversarial perturbations into the diffusion training process, thereby strengthening the models’ ability to purify adversarial perturbations. Additionally, we propose Rank-Based Gaussian Mapping (RBGM) to improve the compatibility of perturbations with diffusion models. Experimental results validate the effectiveness of ADDT. In conclusion, our study suggests that future research on DBP can benefit from a clearer distinction between stochasticity-driven and purification-driven robustness.
@@ 1 Introduction
Figure 1: Comparison of attack trajectories under different evaluation settings. The attack trajectory in the standard white-box setting significantly deviates from the DW-box trajectory and demonstrates lower effectiveness.
@@
Deep learning has achieved remarkable success across various domains, including computer vision (He et al., 2016), natural language processing (OpenAI, 2023), and speech recognition (Radford et al., 2023). However, within this flourishing landscape, the persistent specter of adversarial attacks casts a shadow over the reliability of these neural models. For vision models, adversarial attacks typically involve introducing imperceptible perturbations into input images, tricking the models into producing incorrect outputs with high confidence (Goodfellow et al., 2015; Szegedy et al., 2014). This vulnerability has spurred substantial research into adversarial defenses (Zhang et al., 2019; Samangouei et al., 2018; Shafahi et al., 2019; Wang et al., 2023).
@@
Recently, diffusion-based purification (DBP) (Nie et al., 2022) has emerged as a promising defense against adversarial attacks. Existing studies suggest that DBP robustness primarily stems from the forward diffusion process, which reduces the distribution gap between clean and adversarial images by applying Gaussian noise (Nie et al., 2022; Wang et al., 2022). While this reduction is theoretically supported, its contribution to DBP robustness has not been sufficiently validated through empirical studies. Additionally, experimental results suggest that the stochastic nature of DBP might also play a significant role in enhancing its robustness (Nie et al., 2022).
@@
In this paper, we present a new perspective that emphasizes the role of stochasticity throughout the DBP process as a key contributor to its robustness, challenging the traditional focus on the forward diffusion process. To assess the impact of stochasticity, we propose a Deterministic White-box (DW-box) attack setting, in which the attacker has complete knowledge of both the model parameters and the stochastic elements. Our findings reveal that DBP models experience a significant loss of robustness when the process is made entirely deterministic from the attacker’s perspective. Further analysis of attack trajectories and the loss landscape shows that DBP models do not defend against adversarial perturbations by relying on a flat loss landscape, as is common in adversarial training (AT) (Madry et al., 2018); instead, they leverage stochasticity to bypass the most effective attack directions, as illustrated in Figure 1.
@@
Building on this new perspective of DBP robustness, we hypothesize that strengthening the diffusion model’s ability to purify adversarial perturbations could further improve the robustness. To test this hypothesis, we propose Adversarial Denoising Diffusion Training (ADDT) for DBP models. This method follows an iterative two-step process: first, the Classifier-Guided Perturbation Optimization (CGPO) step generates adversarial perturbations; then, the diffusion model training step updates the parameters of the diffusion model using these perturbations. To better integrate these perturbations into the diffusion framework, we propose Rank-Based Gaussian Mapping (RBGM), which adjusts the adversarial perturbations to more closely resemble Gaussian noise, in line with the theoretical foundation of diffusion models. Experiments confirm that ADDT consistently enhances the robustness of DBP models. Through further empirical analysis and discussion, we argue that future research on DBP should separate the robustness derived from stochasticity and that achieved through purification. This distinction points to two complementary directions for improving DBP: (1) enhancing its ability to purify adversarial perturbations through efficient training methods, and (2) defending against Expectation over Transformation (EoT) attacks (Athalye et al., 2018b) by increasing the variance of attack gradients.
@@
  • We offer a novel perspective on DBP robustness, highlighting the crucial role of stochasticity while challenging the conventional, purification-based view that robustness primarily arises from distribution gap reduction during the forward diffusion process.
  • We introduce a new Deterministic White-box attack setting and demonstrate that DBP models rely on stochastic attack gradients to evade the most effective attack directions, revealing key differences from adversarially trained models.
  • We show that DBP robustness can be further improved by enhancing the diffusion model’s ability to purify adversarial perturbations through the proposed ADDT method.
@@ 2 Related Work
Adversarial training (AT). First introduced by Madry et al. (2018), AT seeks to develop a robust classifier by incorporating adversarial examples into the training process. It has nearly become the de facto standard for improving the adversarial robustness of neural networks (Athalye et al., 2018a; Gowal et al., 2020; Rebuffi et al., 2021). Recent advances in AT harness the generative power of diffusion models to augment training data and prevent overfitting (Gowal et al., 2021; Wang et al., 2023). However, the application of AT to DBP methods has not been thoroughly explored.
Adversarial purification. Adversarial purification utilizes generative models to remove adversarial perturbation from inputs before they are processed by downstream models. Traditionally, generative adversarial networks (GANs) (Samangouei et al., 2018) or autoregressive models (Song et al., 2018) have been used as purification models. More recently, diffusion models have been introduced for adversarial purification in a technique known as diffusion-based purification (DBP), which has shown promising results (Nie et al., 2022; Wang et al., 2022; Wu et al., 2022; Xiao et al., 2022). The robustness of DBP models is often attributed to the wash-out effect of Gaussian noise introduced during the forward diffusion process. Nie et al. (2022) propose that the forward diffusion process reduces the Kullback-Leibler (KL) divergence between the distributions of clean and adversarial images. Gao et al. (2022) suggest that while the forward diffusion process improves robustness by reducing model invariance, the backward process restores this invariance, thereby undermining robustness. However, these theoretical explanations lack strong experimental support.
@@ 3 Preliminaries
Adversarial training. Adversarial training aims to build a robust model by including adversarial samples during training (Madry et al., 2018). This approach can be formulated as a min-max problem, where it first generates adversarial samples (the maximization step) and then adjusts the parameters to resist these adversarial samples (the minimization step). Formally, this can be represented as:
$\bm{\theta}^{*}={\operatorname{arg\,min}}_{\bm{\theta}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[{\max}_{\bm{\delta}\in\mathcal{B}}L(f(\bm{\theta},\bm{x}+\bm{\delta}),y)\right],$ (1)
where $L$ is the loss function, $f$ is the classifier, $(\bm{x},y)\sim\mathcal{D}$ denotes the sampling of training data from the distribution $\mathcal{D}$, and $\mathcal{B}$ defines the set of permissible perturbation $\bm{\delta}$.
Diffusion models. Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) and Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) simulate a gradual transformation in which noise is progressively added to an image and then removed to restore the original image. The forward process can be represented as:
$\bm{x}_{t}=\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (2)
where $\bm{x}_{0}$ is the original image and $\bm{x}_{t}$ is the noisy image. $\overline{\alpha}_{t}$ is the cumulative noise level at step $t$ ($1<t\leq T$, where $T$ is the number of diffusion training steps). The parameters $\bm{\theta}$ are optimized by minimizing the distance between the actual and estimated noise:
$\bm{\theta}^{*}={\operatorname{arg\,min}}_{\bm{\theta}}\mathbb{E}_{\bm{x}_{0},t,\bm{\epsilon}}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\sqrt{\overline{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\bm{\epsilon},t)\|_{2}^{2}\right],$ (3)
where $\bm{\theta}^{*}$ represents the optimized parameters, and $\bm{\epsilon}_{\bm{\theta}}$ is the model’s predicted noise. Using $\bm{\epsilon}_{\bm{\theta}^{*}}$, we can estimate $\hat{\bm{x}}_{0}$ in a single step:
@@
where $\beta_{t}$ is the noise level at step $t$, $\hat{\bm{x}}_{t-1}$ is the image recovered at step $t-1$, and $\bm{\epsilon}$ is sampled from $\mathcal{N}(\bm{0},\bm{I})$. DDIM speeds up the denoising process by skipping certain intermediate steps. Recent work suggests that DDPM could also benefit from a similar approach (Nichol & Dhariwal, 2021). Score SDEs (Song et al., 2021a) provide a score function perspective on DDPM and further lead to the derivations of DDPM++ (VPSDE) and EDM  (Karras et al., 2022). In this diffusion process, the noise terms $\bm{\epsilon}$ in Equation 2 and Equation 5 represent the key stochastic elements that control the randomness of the process. More discussion of stochastic elements is provided in Section C.1.
@@ 4.1 Stochasticity as the Main Factor of DBP Robustness
As discussed in Section 2, previous studies primarily attribute the robustness of DBP models to the forward diffusion process, which perturbs inputs with Gaussian noise to reduce the distribution gap between adversarial and clean images (Nie et al., 2022; Wang et al., 2022). As a result, adversarial perturbations can be “washed out” by the Gaussian noise. However, it has also been observed that the robustness of DiffPure diminishes when switching from Stochastic Differential Equation (SDE) sampling to Ordinary Differential Equation (ODE) (Nie et al., 2022), which introduces less stochasticity. This reduction in robustness cannot be fully explained by the “wash out” effect of Gaussian noise, suggesting that stochasticity plays a role in DBP robustness.
To assess the impact of stochasticity on DBP robustness, we implemented both DDPM and DDIM within the DiffPure framework (Nie et al., 2022) and compared their performance (referred to as DPDDPM and DPDDIM). Note that the original implementation of DiffPure adopts a DDPM discretization form of DDPM++ (VPSDE), which differs only minimally from DDPM (Ho et al., 2020). Thus, the primary difference between DiffPure and our DPDDPM is that DiffPure employs a larger UNet. DDPM utilizes a stochastic SDE-based reverse process, which introduces Gaussian noise in both the forward and reverse processes, making the entire process stochastic. In contrast, DDIM uses a deterministic ODE-based reverse process and introduces Gaussian noise only in the forward process. The results are shown in Figure 2 (labeled as Clean and White), where White refers to $\ell_{\infty}$ white-box PGD (Madry et al., 2018) + EoT (Athalye et al., 2018b) attacks (as detailed in Section 6.1) Although DPDDPM achieves higher clean accuracy, it shows lower robust accuracy under adaptive white-box attacks, consistent with the findings of Nie et al. (2022). This suggests that the stochasticity in the reverse diffusion process may also play a crucial role in DBP robustness.
@@
Figure 2: DPDDPM and DPDDIM robust accuracy under different attack settings on CIFAR-10. Both models lose most of their robustness only when the attacker knows all stochastic elements (highlighted in bold: DWBoth-box for DPDDPM/DPDDIM and DWFwd-box for DPDDIM).
To better control the stochasticity in both the forward and reverse diffusion processes and to isolate its effect on DBP robustness, we introduce a novel attack setting called the Deterministic White-Box (DW-box) setting. In this setting, the attacker not only has complete knowledge of the model parameters but also of the exact values of the stochastic elements sampled during evaluation, effectively making the diffusion process deterministic from the attacker’s perspective. This setting could be realistic if the attacker knows the seed or the initial random state used for pseudo-random number generation in the model. For our evaluations, we define three levels of attacker knowledge: (1) Conventional White-box setting, where the attacker has access to the model parameters but not the stochastic elements; (2) DWFwd-box/DWRev-box setting, where the attacker knows the stochastic elements in the forward/reverse process in addition to the model parameters; (3) DWBoth-box setting, where the attacker has complete knowledge of the model parameters and all stochastic elements in both the forward and reverse processes. Details of these settings are provided in Section C.2.
With the Deterministic White-Box attack, we are able to compare traditional theories with our proposed hypothesis. These explanations diverge in behavior when stochasticity is controlled. Traditional theories emphasize the forward diffusion process as the primary defense mechanism, suggesting that both DPDDPM and DPDDIM should behave similarly under the DWFwd-box setting. In contrast, our hypothesis emphasizes the stochasticity throughout the diffusion process as the crucial factor. We hypothesize that as DPDDIM becomes deterministic under the DWFwd-box setting, it should experience a significant reduction in robustness—similar to DPDDPM under the DWBoth-box setting. We evaluated adversarial robustness on CIFAR-10 using $\ell_{\infty}$ attacks (see Section 6.1). The results are shown in Figure 2. Under the DWFwd-box setting, DPDDPM maintains a significant portion of its robustness, whereas DPDDIM loses almost all resistance to adversarial attacks. This phenomenon can not be explained by the “wash out” theory. Furthermore, under the DWBoth-box setting, DPDDPM shows a substantial drop similar to DPDDIM in the DWFwd setting. This suggests that stochasticity in both the forward and reverse diffusion processes plays a critical role in maintaining robustness.
Our findings suggest that DBP models primarily rely on stochasticity to resist adversarial attacks, rather than depending solely on the forward diffusion process, and it also reveals that DBP models lack the ability to effectively purify adversarial perturbations.